OpenStudy (n8kd_da_beast):

cos (-theta)=(sqrt 7)/2 sin theta <0 what is the value of sin theta? A. -(sqrt 2)/3 B. -(2 sqrt 7)/3 C. -(sqrt 14)/2 D. (sqrt 14)/7 Best answers get a medal!!

rishavraj (rishavraj):

u see \[\cos (-\theta) = \cos \theta ~~~and~~\sin(- \theta) = -\sin \theta\]

OpenStudy (n8kd_da_beast):

ok

OpenStudy (n8kd_da_beast):

@Lexi.babyy

OpenStudy (n8kd_da_beast):

@Nnesha

OpenStudy (n8kd_da_beast):

@FakeReality